OpenStudy (anonymous):

darnell is selecting names from a hat to determine who will choose the family activities for the weekend. The names in the hat include his mother, father, two sister, and him. Calculate the probability of each compound event. Term: P(Darnell and sister), without replacement P(mother and father), with replacement P(eithersister and eithersister), with replacement P(parent andsister), without replacement Definition: A) 10% B) 4% C) 16% D) 20%

OpenStudy (anonymous):

5 names in the hat. P(D and a sister) without replacement would be 1/5 x 2/4 = 2/20 = 1/10 since his name is one out of 5, then there are 2 sisters in the next draw (with 4 total). Use same method for other 3 questions...

OpenStudy (anonymous):

\[P(D,\S)=1/5*1/2=1/10=A\] \[P(M,F)=1/5*1/5=1/25=B\] \[P(\S,\S)=2/5*2/5=4/25=C\] \[P(P,\S)=2/5*2/4=1/5=D\]